OpenStudy (alexwee123):

find the derivative of y=x^x in two different ways i got 1 way down by making x^x=xx = ex*ln(x) and just taking it from there...

OpenStudy (lgbasallote):

...honestly idk what you did haha

OpenStudy (lgbasallote):

but here's one way (idk if this is what you did) let y = x^x ln both sides \[\ln y = x \ln x\]implicitly differentiate \[\frac 1y y' = 1 + \ln x\] corss multiply\[y' = y(1 + \ln x)\] \[y' = x^x (1 + \ln x)\]

OpenStudy (anonymous):

another way: let x^x= y take natural log on both sides x (ln x) = ln y Now, differentiate both sides w.r.t. x [use product rule for LHS and for RHS use chain rule] I hope that makes sense..

OpenStudy (alexwee123):

ya i thought of differentiation too :o

OpenStudy (anonymous):

@lgbasallote hit the nail on the head... That is it..

OpenStudy (lgbasallote):

is that what you did @alexwee123 ?

OpenStudy (alexwee123):

uh ya kinda but i got confused for some reason and never rlly did it :/but now i get it :D

OpenStudy (anonymous):

No, @alexwee123 did is: x^x = e^ ln(x^x) = e^(x lnx) Now, he differentiated using the long chain rule

OpenStudy (lgbasallote):

hmm i didnt know about that :o

OpenStudy (lgbasallote):

that looks fun

OpenStudy (anonymous):

Suggestion Please!!! I am studying the Multivariable Calculus from MIT Open Courseware lectures, yes the course 18.02.. and I feel like not understanding the topics fully. can anyone suggest me some better course on Multivariable Calculus. I googled an d saw that UC Berkeley has a course on that. Is that better than that of MIT?

OpenStudy (anonymous):

@lgbasallote I would appreciate a response from you.. for some reason system isn't letting me send you a message...

OpenStudy (lgbasallote):

really? it should....anyway...im not really familiar with websites...im very naive in the net hah...bottomline idk any website

OpenStudy (lgbasallote):

or any US school

OpenStudy (anonymous):

Okay.. no probs